i have a 3 part question that i can't seem to figure out. graph the below two equation --- P =2Q - 2 on vertical and Q=3P -14 on hortizontal, I tried doing this using the y=mx+b but it doesn't seem right.. 2nd Question: find combo of P & Q which meets both constraints and then finally find the slope of each line (looking for derivative of P with respect to Q.) I have tried several options and my answer does not look right. thanks
Answer by stanbon(75887)

Graph the below two equation --- P =2Q - 2 on vertical and Q=3P -14 on hortizontal, I tried doing this using the y=mx+b but it doesn't seem right..
P=2Q-2
P=(Q+14)/3


2nd Question: find combo of P & Q which meets both constraints(what constraints?).
3rd Find the slope of each line (looking for derivative of P with respect to Q.)
Derivative of P=2Q-2 is "2"
Derivative of P=(q+14)/3 is (1/3)
Cheers,
Stan H.